OpenStudy (anonymous):

line AB contains points (2,1) and (-1,-8). What is the equation of the line parallel to line AB that contains point (0,2)?

OpenStudy (anonymous):

First, get the slope of line AB:\[\frac{ y _{1}-y _{2} }{ x _{1}-x _{2} } = m\] where "m" is the slope and (x1, y1), and (x2, y2) are the points given in AB.

OpenStudy (anonymous):

Next, the equation of the parallel line is:\[y - y _{1} = m(x - x _{1})\]where (x1, y1) is the point (0, 2). This is the point-slope equation of the line. The line can be put into other forms from this equation, including the standard form and the slope-intercept form.

OpenStudy (anonymous):

i just plug those into the equation

OpenStudy (anonymous):

Well, there are 2 equations, one for each of my previous posts. You have to start with getting "m" which is the slope. Yes, make substitutions.

OpenStudy (anonymous):

ok, i got it now. now that I have seen the equaions. thanks

OpenStudy (anonymous):

yw!